OpenStudy (blaque678):

Graph the line with slope 2,passing through the point (-3,2)

OpenStudy (nubeer):

first u have to find the equation of the line.. can u find the equation fo line?

OpenStudy (blaque678):

y = mx + b

OpenStudy (nubeer):

tell me what equation u got?

OpenStudy (nubeer):

ok i am assuming u did right.. so u got equation so first put x = 0 u will get value of y and put it on graph. now put y = 0 u will get the value of x and plot it on graph.. now u have 2 points on graph join those 2 points that will be ur graph

OpenStudy (blaque678):

y=2x+8

OpenStudy (nubeer):

ok i am assuming u did right.. so u got equation so first put x = 0 u will get value of y and put it on graph. now put y = 0 u will get the value of x and plot it on graph.. now u have 2 points on graph join those 2 points that will be ur graph

OpenStudy (blaque678):

thats where I get mixed up at when it comes to graphing

OpenStudy (nubeer):

first put x = 0 in ur equation.. what value of y u will get?

OpenStudy (nubeer):

nooo.. y = 2(0) +8 y = ?

OpenStudy (blaque678):

8

OpenStudy (nubeer):

yes.. now plot this point on graph..

OpenStudy (blaque678):

im really guessing your miss understanding my question.I have no problem getting the numbers my problem is putting them on the graph.It doesnt come out correct

OpenStudy (nubeer):

 this is ur line

OpenStudy (nubeer):

oops sorry..give me a moment i get ur Q now..

OpenStudy (nubeer):

no its right bro.. this is ur graph..

OpenStudy (nubeer):